### v4

# Fuzzy Classification Aggregation

Abstract
We consider the problem where a set of individuals has to classify $m$ objects into $p$ categories and does so by aggregating the individual classifications. We show that if $m\geq 3$, $m\geq p\geq 2$, and classifications are fuzzy, that is, objects belong to a category to a certain degree, then an optimal and independent aggregator rule that satisfies a weak unanimity condition belongs to the family of Weighted Arithmetic Means. We also obtain characterization results for $m=p=2$.
Keywords: Classification Aggregation; Weighted Arithmetic Mean; Fuzzy Setting.
JEL Classification: D71.
1 Introduction
The study of the problem of individuals classifying objects can be traced back to Kasher and Rubinstein, (1997). In their paper, they consider a finite society that has to determine which one of its subsets of members consists of exactly those individuals that can be deemed to be part of a group named $J$. They propose different sets of axioms characterizing three aggregators: the liberal one, where each individual decides by herself whether she belongs or not to the group; the oligarchic one, where a subset of individuals defines who belongs to the group; and the dictatorial, where a unique individual decides who is a $J$. Kasher and Rubinstein, (1997) gave rise to what can be regarded as a subdomain within social choice theory, namely, the study of the Group Identification Problem (see, among others, Samet and Schmeidler,, 2003; Miller,, 2008; Fioravanti and Tohmé,, 2021, for more details).
A more general problem, where a group of individuals has to classify $m$ objects into $p$ different categories, has been considered by Maniquet and Mongin, (2016). In their work, they study the case where there are at least as many objects as categories, at least three categories and all the categories must be filled with at least one object. [^2] This surjectivity condition becomes relevant, for example, in situations where $m$ workers have to be assigned to $p$ tasks, and no task can be left unassigned. It can also be seen as a sanity check of the classification process. For instance, consider a set of neural networks that is being trained to classify images into images of cats, dogs and rabbits. If a database with a large number of images of each of the three animals is used, then a neural network that fails to classify at least one image as a rabbit can be considered as not reliable. This aspect is combined with three other seemingly natural properties, namely Unanimity, which states that when all individuals make the same classification then the aggregate classification complies; Independence, which is the requirement that an object be classified by the aggregator in the same way at any two classification profiles if every individual classifies it equally in both profiles; and Non-Dictatorship, which requires that there is no individual such that her classification is always selected. The result is that there is no aggregator satisfying these three conditions. Recently, Cailloux et al., (2024) weaken the Unanimity axiom and find a weakening of the impossibility result that holds for $m>p\geq 2$, with the existence of an essential dictator. This is an individual such that a permutation of her classification is always selected. They also characterize the unanimous and independent aggregators when $m=p=2$.
The surjectivity condition in Maniquet and Mongin, (2016), turns out to be a rather demanding requirement. Without it, the impossibility result no longer holds, and many non-dictatorial aggregators satisfying Unanimity and Independence can be found in the literature (See, for example, Kasher and Rubinstein,, 1997).
Our work considers the problem where $m$ objects have to be classified into $p$ categories, $m\geq p\geq 2$, and the classifications, both by the individuals and the rule, indicate degrees of membership of the objects to each of the categories. Two possible scenarios where this setting may arise, are the following. A national government has to allocate equal funds to all regions, to be used for different regional ministries (security, education, and health), with the central government providing some conditions on how the funds are distributed. It is natural to assume in this context, that the government sets a minimum amount of money that should be spent in each of the ministries, at least at a national level. Members of the government should decide on how the funds might be allocated for security, education, and health, ensuring that each category, across the country, receives at least the amount provided by the government to each region. As an illustration, one region could be tasked with allocating half of the funds to security, a quarter to education, and a quarter to health. Another scenario involves assigning individuals to various mandatory tasks with equal time requirements, where proportions dictate how each person’s time is allocated. For instance, if tasks include laundry, lawn mowing, and grocery shopping, Bob might spend half of his time on laundry, a quarter on mowing the lawn, and the remaining quarter on grocery shopping. [^3] These examples illustrate why the notion of “degree of membership” can be considered natural to use, instead of considering probabilities. For instance, in the tasks allocation example, it is better if we know that Bob will spend with certainty half of his time doing the laundry, rather than knowing that there is a $50\%$ chance of Bob leaving all the clothes dirty.
Fuzzy preferences, those that represent vagueness and uncertainty, are a useful tool that has been used in many aggregation problems. Noteworthy contributions include the work of Dutta, (1987), who deals with exact choices under vague preferences, Dutta et al., (1986) who investigate the structure of fuzzy aggregation rules determining fuzzy social orderings, and recent work by Duddy and Piggins, (2018), who prove the fuzzy counterpart of Weymark’s general oligarchy theorem (Weymark,, 1984), and by Raventos-Pujol et al., (2020), who analyze Arrow,’s (1951) theorem in a fuzzy setting.
Numerous studies have addressed the Group Identification Problem when the preferences or classifications are not crisp. For instance, Cho and Park, (2018) present a model of group identification for more than two groups, allowing fractional classifications but no fractional opinions, Ballester and García-Lapresta, (2008) deal with fuzzy opinions in a sequential model, and Fioravanti and Tohmé, (2022) show that some of the impossibility results of Kasher and Rubinstein, (1997) can be avoided. Alcantud et al., (2019) analyze the classification problem in a fuzzy setting, and consider a strong fuzzy counterpart of the surjectivity condition, extending the impossibility result of Maniquet and Mongin, (2016).
We avoid both impossibility results, the crisp and the fuzzy ones, by considering a different, but also very natural, fuzzy surjectivity condition. This condition, which in the crisp setting is extremely demanding, becomes a requirement that in the fuzzy setting allows the existence of a family of aggregators satisfying certain particular properties. Thus, it is demanding enough to find a characterization result, but not so much to not be able to find one. Building on some previous results in functional analysis by Aczél and Wagner, (1980); Aczél et al., (1984), and Wagner, (1982), we show that the only optimal aggregators that are independent and that satisfy a weak version of Unanimity, are the Weighted Arithmetic Means.
Considering different settings for already established problems is a common technique used in social choice to avoid impossibility results. The result by Gibbard, (1973) and Satterthwaite, (1975), that states in the deterministic setting that the only rule immune to manipulations is the dictatorial one, is avoided when the outcomes of the rule are lotteries over the set of alternatives (Gibbard,, 1977). We consider that our paper is related to Maniquet and Mongin, (2016)’s paper in an analogous way as Dutta, (1987) and Billot, (1991) results are related to the Arrow,’s (1951) impossibility theorem. In their work, they consider a fuzzy setting and adapt Arrow’s theorems to it to obtain possibility results. Dutta, (1987) characterizes the oligarchic aggregators, while Billot, (1991) does the same for the leximin aggregator. The use of a fuzzy setting, that allows many interpretations of axioms that are used in crisp settings, proves advantageous for circumventing impossibility results.
The plan of the paper is as follows. Section 2 presents the basic notions and axioms that we use, while we present the results in Section 3. Finally, Section 4 contains some concluding remarks.
2 Basic Notions and Axioms
Let $N=\{1,\ldots,n\}$ be a finite set of individuals and let $X=\{x_{1},\ldots,x_{m}\}$ be $m$ objects that need to be classified into the $p$ categories of a set $P$, with $p\geq 2$.
The individuals classify each object according to a partial degree of membership to each category. In the crisp setting, introduced by Maniquet and Mongin, (2016), classifications are surjective mappings $c:X\rightarrow P$, that is, every category must have at least one object classified into. In the fuzzy setting, a fuzzy classification is a mapping $c:X\rightarrow[0,1]^{P}$ such that $\sum_{j=1}^{m}c(x_{j})_{t}\geq 1$ for each $t\in\{1,\ldots,p\}$ and $\sum_{t=1}^{p}c(x)_{t}=1$ for all $x\in X$. The former condition is the fuzzy counterpart of the surjectivity of the classification function, while the latter is the fuzzy counterpart of the assumption that every object $x$ must be assigned to exactly one category by each voter. It is easy to see that these conditions imply that there must be at least as many objects as categories, thus $m\geq p$.
We use $\mathcal{C}$ to denote the set of fuzzy classifications, and every $\bm{c}=(c_{1},\ldots,c_{n})\in\mathcal{C}^{N}$ is a fuzzy classification profile. Given $\mathbf{c}\in\mathcal{C}^{N}$ and $x\in X$, we denote with $\mathbf{c}^{x}\in\mathcal{C}^{N}|_{x}$ the fuzzy classification profile restricted to x, such that the entry $\mathbf{c}^{x}_{ij}$ indicates the degree of membership of the object $x$ to the category $j$, according to the agent $i$. Thus, we have that $\bm{c}^{x}_{ij}=c_{i}(x)_{j}$. This allows the representation of a fuzzy classification profile as a set of $m$ matrices in a subset of $[0,1]^{N\times P}$, one for each object, where row $i$ indicates the classifications given by the individual $i$, and column $t$ of the matrix indicates the classifications given by the set of individuals into the category $t$.
A fuzzy classification aggregation function (FCAF) is a mapping $\alpha:\mathcal{C}^{N}\rightarrow\mathcal{C}$ such that $\alpha(\bm{c})(x)$ indicates the degrees of membership to the different categories of the object $x$. We call the outcome of $\alpha$, the fuzzy social classification.
Next, we introduce a particular FCAF, the Weighted Arithmetic Mean. Let $\bm{w}=(w_{1},\ldots,w_{n})$ be a set of weights such that $w_{i}\in[0,1]$ for all $i\in N$ and $\sum_{i=1}^{n}w_{i}=1$. Then $\alpha_{\bm{w}}:\mathcal{C}^{N}\rightarrow\mathcal{C}$ is such that for all $x\in X$,
$\alpha_{\bm{w}}(\bm{c})(x)=w_{1}c_{1}(x)+\cdots+w_{n}c_{n}(x)$
We say that the set of weights is degenerate if there is an $i\in N$ such that $w_{i}=1$. If we think of the FCAF as a group of experts classifying objects, this FCAF can be appropriate for situations where the individuals differ in expertise, with more experienced voters having higher weights. [^4] Weighted aggregators have been previously studied in the social choice literature, in particular in the setting where monetary transfers are allowed (see Roberts,, 1979; Mishra and Sen,, 2012). See also Stone, (1961) and Dietrich and List, (2016) for an use of this aggregator in opinion pooling. In particular, if $w_{i}=\frac{1}{n}$ for all $i\in N$, we call this FCAF the Arithmetic Mean.
Example 1.
Consider a situation where there are $n=3$ individuals that have to classify $m=3$ objects into $p=3$ categories. Let $\bm{c}$ be a fuzzy classification profile such that:
$\bm{c}^{x_{1}}=\begin{pmatrix}\frac{1}{3}&\frac{2}{3}&0\\ 1&0&0\\ \frac{1}{2}&\frac{1}{4}&\frac{1}{4}\\ \end{pmatrix}$, $\bm{c}^{x_{2}}=\begin{pmatrix}\frac{2}{3}&0&\frac{1}{3}\\ 0&0&1\\ \frac{1}{2}&0&\frac{1}{2}\\ \end{pmatrix}$, and $\bm{c}^{x_{3}}=\begin{pmatrix}0&\frac{1}{3}&\frac{2}{3}\\ 0&1&0\\ 0&\frac{3}{4}&\frac{1}{4}\\ \end{pmatrix}$.
In this example, individual $2$ classifies the object $x_{2}$ as being only part of category $3$ (second row of the second matrix). If we consider the Weighted Arithmetic Mean with a non-degenerate set of weights $\bm{w}=(\frac{1}{2},0,\frac{1}{2})$, we obtain the following fuzzy social classification:
$\alpha_{\bm{w}}(\bm{c})(x_{1})=(\frac{10}{24},\frac{11}{24},\frac{3}{24})$, $\alpha_{\bm{w}}(\bm{c})(x_{2})=(\frac{14}{24},0,\frac{10}{24})$, and $\alpha_{\bm{w}}(\bm{c})(x_{3})=(0,\frac{13}{24},\frac{11}{24})$.
In the following, we introduce a number of axioms, i.e., fundamental normative requirements that, depending on the specific classification problem at hand, a reasonable FCAF should satisfy. The first axiom can be seen as an efficiency property. It states that if all the individuals agree on how many objects should be classified into a given category, then the social classification must respect that.
Definition 1 (Optimality).
An FCAF is optimal if for all $\bm{c}\in\mathcal{C}^{N}$ such that $\sum_{j=1}^{m}c_{i}(x_{j})_{t}=h$ for all $i\in N$, it is the case that $\sum_{j=1}^{m}\alpha(\bm{c})(x_{j})_{t}=h$.
This axiom is useful in a setting where categories need to fulfil certain requirements and resources should not be wasted. For example, in the case where a set of managers has to classify a group of workers as members of different departments within a company, in order to make them operative. If all the managers consider that department A requires only two workers to be operative, then an optimal aggregator only classifies two workers into that department. The following axiom states that the fuzzy social classification of an object in two different fuzzy classification profiles does not change if the classification regarding that object is the same in both profiles for every individual.
Definition 2 (Independence).
An FCAF is independent if given $\bm{c,c^{\prime}}\in\mathcal{C}^{N}$ and $x\in X$ such that $c_{i}(x)=c^{\prime}_{i}(x)$ for all $i\in N$, it is the case that $\alpha(\bm{c})(x)=\alpha(\bm{c^{\prime}})(x)$.
An independent FCAF $\alpha$ can be seen as a collection of mappings $(\alpha_{x})_{x\in X}$, such that $\alpha_{x}:\mathcal{C}^{N}|_{x}\rightarrow\mathcal{C}|_{x}$ and $\alpha_{x}(\bm{c}^{x})=\alpha(\bm{c})(x)$ (Cailloux et al.,, 2024). We call these mappings Elementary FCAFs. A stronger version of Independence requires that if two objects are classified equally in a profile, then the fuzzy social classification must be the same. [^5] This property is called Strong Label Neutrality in Wagner, (1982).
Definition 3 (Symmetry).
An FCAF is symmetric if for all $x,y\in X$ and all $\bm{c}\in\mathcal{C}^{N}$ such that $c_{i}(x)=c^{\prime}_{i}(y)$ for all $i\in N$, it is the case that $\alpha(\bm{c})(x)=\alpha(\bm{c})(y)$.
Wagner, (1982) shows that an optimal and symmetric FCAF is also independent, and thus has the same elementary FCAF for every object. The next property states that if there is an object that is unanimously classified by the individuals, then the FCAF has to classify that object accordingly.
Definition 4 (Unanimity).
An FCAF is unanimous if for all $\bm{c}\in\mathcal{C}^{N}$ where there is an $x\in X$ and a category $t\in P$ such that $c_{1}(x)_{t}=\cdots=c_{n}(x)_{t}=r$, it is the case that $\alpha(\bm{c})(x)_{t}=r$.
A weaker version of Unanimity states that the only unanimous classification required to be respected is when all the individuals classify an object with a degree of membership 00. The intuition behind this axiom is that if all individuals consider that an object does not belong at all to a certain category, then the social classification should agree.
Definition 5 (Zero Unanimity).
An FCAF is zero unanimous if for all $\bm{c}\in\mathcal{C}^{N}$ where there is an $x\in X$ and a category $t\in P$ such that $c_{1}(x)_{t}=\cdots=c_{n}(x)_{t}=0$, it is the case that $\alpha(\bm{c})(x)_{t}=0$.
Now we introduce our first fuzzy axiom, which can be seen as the fuzzy counterpart of Unanimity. [^6] This axiom is called coherence by Alcantud et al., (2019). Here we use the same name used by Fioravanti and Tohmé, (2022). It states that the degree to which an object is collectively classified must be between the classification degrees regarding that object for each of the individuals.
Definition 6 (Fuzzy Consensus).
An FCAF satisfies Fuzzy Consensus if for every object $x\in X$ and for every category $t\in\{1,\ldots,p\}$, it is the case that $\alpha(\mathbf{c})(x)_{t}\in[\min_{i\in N}c_{i}(x)_{t},\max_{i\in N}c_{i}(x)_{t}]$.
It is easy to see that Fuzzy Consensus implies Unanimity, and Unanimity implies Zero Unanimity. These three axioms can be interpreted as the different ‘degrees’ of consensus that we might require an aggregator to satisfy. The next axiom states that there must not exist an individual that imposes her classification.
Definition 7 (Non-Dictatorship).
An FCAF is non-dictatorial if there is no individual $i\in N$ such that for all $x\in X$ and all $\bm{c}\in\mathcal{C}^{N}$, it is the case that $\alpha(\mathbf{c})(x)=c_{i}(x)$.
A stronger version of Non-Dictatorship states that the names of the individuals are not important for the aggregation of classifications.
Definition 8 (Anonymity).
For a classification profile $\bm{c}\in\mathcal{C}^{N}$ and a permutation $\sigma:N\rightarrow N$ we define $\sigma(\bm{c})$ as the classification profile such that $\sigma(\bm{c})=\{\bm{c}_{\sigma(1)},\ldots,\bm{c}_{\sigma(n)}\}$. An FCAF is anonymous if for all $\bm{c}\in\mathcal{C}^{N}$ and all permutations $\sigma:N\rightarrow N$, it is the case that $\alpha(\sigma(\bm{c}))=\alpha(\bm{c})$.
3 Results
In the crisp setting, Maniquet and Mongin, (2016) show the following impossibility result.
Theorem.
(Maniquet and Mongin,, 2016) Let $m\geq p\geq 3$. There is no (non-fuzzy) Classification Aggregation Function that satisfies Independence, Unanimity, and Non-Dictatorship.
Alcantud et al., (2019) extend this result to a fuzzy setting, by using Fuzzy Consensus and considering a stronger notion of surjectivity, where for each category there must exist an object with a degree of classification larger than $0.5$. They show the existence of a fuzzy dictator, where whenever the dictator classifies an object into a category by more than $0.5$, then the object is classified into that category with a degree of more than $0.5$. Thus, our main theorem can be seen as an escape from these impossibility results, even in a fuzzy setting.
Theorem 1.
Let $m\geq 3$ and $m\geq p\geq 2$. An optimal Fuzzy Classification Aggregation Function satisfies Independence, Zero Unanimity, and Non-Dictatorship if, and only if, it is a Weighted Arithmetic Mean with a non-degenerate set of weights.
Now we present a result by Aczél et al., (1984) that is useful for our proof.
Theorem.
(Aczél et al.,, 1984) A family of mappings $\{\alpha^{j}:[0,1]^{N}\rightarrow[0,1]\}_{j=1}^{m}$ satisfies Zero Unanimity and the following condition:
$\text{if }z^{j}\in[0,1]^{N}\text{ and }\sum_{j=1}^{m}z^{j}=(1,\ldots,1),\text{then }\sum_{j=1}^{m}\alpha^{j}(z^{j})=1$ (1)
if, and only if, there exists a set of weights $(w_{1},\ldots,w_{n})$ such that for $j=1,\ldots,m$ and any $z\in[0,1]^{N}$, we have that
$\alpha^{j}(z)=w_{1}z_{1}+\cdots+w_{n}z_{n}.$
Proof of Theorem 1.
That a Weighted Arithmetic Mean with a non-degenerate set of weights satisfies the four axioms is straightforward to see. For the only if part, we use the result by Aczél et al., (1984). Let $\alpha$ be an optimal FCAF that satisfies Independence and Zero Unanimity. Recall that an independent FCAF can be seen as a collection of elementary FCAFs. So we can consider a family of mappings $\{\alpha_{x_{j}}:\mathcal{C}^{N}|_{x_{j}}\rightarrow\mathcal{C}|_{x_{j}}\}_{j=1}^{m}$. By optimality, we obtain that if $\sum_{j=1}^{m}c_{i}(x_{j})_{t}=h$ for all $i\in N$, it is the case that $\sum_{j=1}^{m}\alpha_{x_{j}}(\bm{c}^{x_{j}})_{t}=h$. Then we have that for each $t\in P$ there is a family of mappings $\{\alpha_{x_{j}}|_{t}:\mathcal{C}^{N}|_{x_{j}\land t}\rightarrow[0,1]\}_{j=1}^{m}$ satisfying condition (1). The function $\alpha_{x_{j}}|_{t}$ maps the column $t$ in $\bm{c}^{x_{j}}$ and outputs the degree of classification of object $x_{j}$ into the category $t$. Due to optimality, regardless of the values that are not in column $t$, we can be sure that condition (1) is satisfied by $\alpha_{x_{j}}|_{t}$ and that the degree of classification of the object $x_{j}$ into the category $t$ depends on a weighted aggregator. Thus we have that for each category $t$ there is a set of weights $(w^{t}_{1},\ldots,w^{t}_{n})$ such that for $j=1,\ldots,m$, it is the case that
$\alpha_{x_{j}}|_{t}(\bm{c}^{x_{j}}_{1t},\ldots,\bm{c}^{x_{j}}_{nt})=w^{t}_{1}\bm{c}^{x_{j}}_{1t}+\cdots+w^{t}_{n}\bm{c}^{x_{j}}_{nt}.$
What is left to show is that for every $i\in N$, the weights are the same for every category, meaning that $w_{i}^{t}=w_{i}^{t^{\prime}}$ for all $t,t^{\prime}\in P$. To avoid excessive notation, we show it for the case of $|N|=2$, and $m=p=3$. It is easy to extend it for the general case. Consider the example where $c_{1}(x_{1})=(1,0,0)$, $c_{1}(x_{2})=(0,1,0)$, $c_{1}(x_{3})=(0,0,1)$, $c_{2}(x_{1})=(0,1,0)$, $c_{2}(x_{2})=(0,0,1)$, and $c_{2}(x_{3})=(1,0,0)$. Thus we obtain $\alpha_{x_{1}}(c_{1}(x_{1}),c_{2}(x_{1}))=(w_{1}^{1},w_{2}^{2},0)$, $\alpha_{x_{2}}(c_{1}(x_{2}),c_{2}(x_{2}))=(0,w_{1}^{2},w_{2}^{3})$, and $\alpha_{x_{3}}(c_{1}(x_{3}),c_{2}(x_{3}))=(w_{2}^{1},0,w_{1}^{3})$, where every three dimensional vector adds up to one. Thus we obtain the following system of equations:
$\begin{cases}w_{1}^{1}+w_{2}^{1}=1&(1)\\ w_{1}^{2}+w_{2}^{2}=1&(2)\\ w_{1}^{3}+w_{2}^{3}=1&(3)\\ w_{1}^{1}+w_{2}^{2}=1&(4)\\ w_{1}^{2}+w_{2}^{3}=1&(5)\\ w_{2}^{1}+w_{1}^{3}=1&(6)\end{cases}$
From $(1)$ and $(4)$ we obtain that $11-w_{2}^{1}+w_{2}^{2}=$ 1, thus $w_{2}^{1}=w_{2}^{2}$, and from $(2)$ and $(5)$ we obtain that $11-w_{2}^{2}+w_{2}^{3}=$ 1, thus $w_{2}^{2}=w_{2}^{3}$. It is a similar procedure for the weights of agent $1$. So the FCAF associates to every individual $i\in N$ a weight $w_{i}$ such that $w_{i}\geq 0$ and that $\sum_{i=1}^{n}w_{i}=1$, and as it is non-dictatorial, we have that there is no $i\in N$ such that $w_{i}=1$. Thus the FCAF is a Weighted Arithmetic Mean with a non-degenerate set of weights. ∎
The impossibility result from the crisp setting is obtained when the aggregator is a Weighted Arithmetic Mean and the set of weights is degenerate. As the aggregator must have a crisp outcome, this is the only possible set of weights (except for permutations on the name of the individuals). One notable aspect of this result is that the weight assigned to each individual remains consistent across all categories, with no category receiving a higher weight than the others. An implication of our main result is that if we want all individuals to be considered the same, then the weights must be the same for all individuals.
Corollary 1.
Let $m\geq 3$ and $m\geq p\geq 2$. An optimal Fuzzy Classification Aggregation Function satisfies Independence, Zero Unanimity, and Anonymity if, and only if, it is an Arithmetic Mean.
As the Weighted Arithmetic Means with non-degenerate weights satisfy Unanimity and Fuzzy Consensus, stronger versions of Zero Unanimity, we have the following two corollaries.
Corollary 2.
Let $m\geq 3$ and $m\geq p\geq 2$. An optimal Fuzzy Classification Aggregation Function satisfies Independence, Unanimity, and Non-Dictatorship if, and only if, it is a Weighted Arithmetic Mean with a non-degenerate set of weights.
Corollary 3.
Let $m\geq 3$ and $m\geq p\geq 2$. An optimal Fuzzy Classification Aggregation Function satisfies Independence, Fuzzy Consensus, and Non-Dictatorship if, and only if, it is a Weighted Arithmetic Mean with a non-degenerate set of weights.
Thus, both impossibility results by Maniquet and Mongin, (2016) and Alcantud et al., (2019) are avoided in our setting, even if we use their same ‘consensual’ axioms.
[…]
Corollary 4.
[…]
It is worth mentioning that the previous results are only valid for $m\geq 2$. For the case of $m=2$ objects and $p=2$ categories, Independence is trivially satisfied (due to the surjectivity conditions). This forces us to use a stronger axiom to obtain a characterization.
For optimal, symmetric and zero unanimous FCAFs, we can have FCAFs that are not Weighted Arithmetic Means, as the following result shows.
Proposition 1.
Let $m=p=2$. An optimal and symmetric Fuzzy Classification Aggregation Function $\alpha$ satisfies Zero Unanimity if, and only if, there is a function $h:[-\frac{1}{2},\frac{1}{2}]^{n}\rightarrow[-\frac{1}{2},\frac{1}{2}]$ where
$h(x_{1},\ldots,x_{n})=-h(-x_{1},\ldots,-x_{n}),$
and
$h\left(\dfrac{1}{2},\ldots,\dfrac{1}{2}\right)=\dfrac{1}{2},$
such that for each $\bm{c}\in\mathcal{C}^{N}$,
$\alpha(\bm{c})(x_{t})=\left(h\left(\bm{c}^{x_{t}}_{11}-\frac{1}{2},\ldots,\bm{c}^{x_{t}}_{n1}-\frac{1}{2}\right)+\frac{1}{2},h\left(\bm{c}^{x_{t}}_{12}-\frac{1}{2},\ldots,\bm{c}^{x_{t}}_{n2}-\frac{1}{2}\right)+\frac{1}{2}\right)$
for $t=1,2$.
Before the proof, we present the following result by Wagner, (1982).
Theorem.
(Wagner,, 1982) Let $m=p=2$. An optimal and symmetric mapping $\alpha:[0,1]^{N\times 2}\rightarrow[0,1]^{2}$ satisfies Zero Unanimity if, and only if, there exists a function $h:[-\frac{1}{2},\frac{1}{2}]^{n}\rightarrow[-\frac{1}{2},\frac{1}{2}]$ where
$h(y_{1},\ldots,y_{n})=-h(-y_{1},\ldots,-y_{n}),$
and
$h\left(\dfrac{1}{2},\ldots,\dfrac{1}{2}\right)=\dfrac{1}{2},$
such that for each $\bm{c}\in[0,1]^{N\times 2}$, we have that
$\alpha(\bm{c})=\left(h\left(\bm{c}_{11}-\frac{1}{2},\ldots,\bm{c}_{n1}-\frac{1}{2}\right)+\frac{1}{2},h\left(\bm{c}_{12}-\frac{1}{2},\ldots,\bm{c}_{n2}-\frac{1}{2}\right)+\frac{1}{2}\right).$
Proof of Proposition 1..
That such an FCAF satisfies Zero Unanimity is straightforward to see. For the if part, we use the result by Wagner, (1982). By Optimality and Zero Unanimity, for each category the hypotheses of Wagner, (1982) are satisfied and thus, we have a mapping $h$ satisfying Wagner,’s conditions, namely $h^{1}$ and $h^{2}$. What is left to see is that $h^{1}=h^{2}$. We have the following conditions that must be satisfied by $\alpha(\bm{c})(x_{1})$ and $\alpha(\bm{c})(x_{2})$:
$h^{1}\left(\bm{c}^{x_{1}}_{11}-\frac{1}{2},\ldots,\bm{c}^{x_{1}}_{n1}-\frac{1}{2}\right)+\frac{1}{2}+h^{2}\left(\bm{c}^{x_{1}}_{12}-\frac{1}{2},\ldots,\bm{c}^{x_{1}}_{n2}-\frac{1}{2}\right)+\frac{1}{2}=1,$
and
$h^{1}\left(\bm{c}^{x_{1}}_{11}-\frac{1}{2},\ldots,\bm{c}^{x_{1}}_{n1}-\frac{1}{2}\right)+\frac{1}{2}+h^{1}\left(\bm{c}^{x_{2}}_{11}-\frac{1}{2},\ldots,\bm{c}^{x_{2}}_{n1}-\frac{1}{2}\right)+\frac{1}{2}=1.$
It is easy to see that $\bm{c}^{x_{1}}_{j2}=\bm{c}^{x_{2}}_{j1}$ for all $i=1,\ldots,n$. Thus we obtain that $h^{1}\left(\bm{c}^{x_{2}}_{11}-\frac{1}{2},\ldots,\bm{c}^{x_{2}}_{n1}-\frac{1}{2}\right)=h^{2}(\bm{c}^{x_{2}}_{11}-\frac{1}{2},\ldots,\bm{c}^{x_{2}}_{n1}-\frac{1}{2})$, concluding our proof.
Despite its technical flavour, Proposition 1 has as a particular case the class of Weighted Arithmetic Means, and also any other power mean with an odd exponent. [^7] A power mean with an odd exponent is a function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ such that $f(x_{1},\ldots,x_{n})=(\frac{1}{n}\sum_{i=1}^{n}x_{i}^{m})^{\frac{1}{m}}$ where $m$ is an odd number. Proposition $1$ of Cailloux et al., (2024) is a particular case of Proposition 1, where condition 1 is implied by the complementary condition of their result.
A more general result, valid for $m\geq p\geq 2$ can be attained if we slightly change the surjectivity conditions. For this general case, we consider fuzzy classifications $c^{\ast}:X\rightarrow\mathbb{R}^{P}$, and the surjectivity requirements are such that for all $x\in X$, and for a given $s\in\mathbb{R}$, it is the case that $\sum_{t=1}^{p}c^{\ast}(x)_{t}=s$, and that $\sum_{j=1}^{m}c^{\ast}(x_{j})_{t}\geq s$ if $s\geq 0$ for all $t\in P$, or $\sum_{j=1}^{m}c^{\ast}(x_{j})_{t}<s$ if $s<0$ for all $t\in P$. We can think of this setting as $n$ operators assigning $s$ hours of use of $m$ machines into $p$ different tasks that need $s$ hours to be finished, or $s$ euros that $m$ persons have to spend in $p$ different projects that need at least $s$ euros to be done. We use $\mathcal{C}^{\ast}$ to denote this set of classifications. Thus, the FCAF is a mapping $\alpha:{\mathcal{C}^{\ast}}^{N}\rightarrow\mathcal{C}^{\ast}$ such that $\alpha(\bm{c}^{\ast})(x_{j})$ indicates, for example, the different hours assigned to the different tasks of the machine $x_{j}$.
We have that an optimal FCAF satisfies the $k$-allocation property for any $k=m\geq 2$ (Aczél and Wagner,, 1980), that is, if for any $t\in P$ such that $\sum_{j=1}^{m}c^{\ast}(x_{j})_{t}=s$ for all $i\in N$, it is the case that $\sum_{j=1}^{m}\alpha(\bm{c}^{\ast})(x_{j})_{t}=s$. The $k$-allocation property allows us to characterize the optimal and symmetric FCAFs that satisfy Fuzzy Consensus for any value of $m$ such that $m\geq p\geq 2$.
Theorem 2.
Let $m\geq p\geq 2$. An optimal and symmetric FCAF satisfies Fuzzy Consensus and Non-Dictatorship if, and only if, it is a Weighted Arithmetic Mean with a non-degenerate set of weights.
Proof.
Aczél and Wagner, (1980) show that a mapping $\bm{c}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ satisfies the k-allocation property for $k=2,3$ and is bounded if, and only if, it is a Weighted Arithmetic Mean. In our setting, symmetry implies that for a given category $t\in P$, we have the same elementary FCAF for every object. By optimality, the k-allocation property is satisfied by the elementary FCAFs. Fuzzy Consensus implies they are bounded and thus they are Weighted Arithmetic Means, and finally, Non-Dictatorship implies that the set of weights is non-degenerate. By a similar proof to the one used in Theorem 1, it is easy to see that the elementary FCAFs are the same for every category, concluding our proof. ∎
4 Final Remarks
In this work, we present an analysis from a fuzzy point of view, of the challenge of classifying $m$ objects into $p$ different categories. The classifications of the objects by the agents and the rules are no longer crisp statements about to which category the object is assigned. Instead of that, the classifications are expressed in terms of degrees of assignment to each of the categories. In the crisp setting for more than two objects and two categories, requiring the rule to fill each category with at least one object, to be independent and unanimous, implies the existence of an agent such that objects are classified according to the opinions of that agent (Maniquet and Mongin,, 2016). Even with a weaker unanimity condition, this result can not have a major improvement (Cailloux et al.,, 2024, showing the existence of an essential dictator).
The fuzzy setting proves advantageous, as the surjectivity conditions can assume diverse yet natural interpretations. Strong interpretations, akin to those in Alcantud et al., (2019), can extend the scope of impossibility results. However, by considering weaker surjectivity conditions and different versions of Unanimity, we can circumvent these limitations. Our findings show that rules that belong to the family of Weighted Arithmetic Means are the only ones that satisfy Optimality, Independence, and Zero Unanimity. Under different interpretations of the classification process and the consensual axioms, we can obtain characterization results that hold for $m\geq p\geq 2$.
An intriguing avenue for further inquiry lies in extending our analysis to dynamic scenarios, where object classifications may evolve over time. Exploring the temporal dynamics of fuzzy classifications could provide valuable insights into the adaptability and stability of the proposed framework in real-world applications.
Declarations of interest: None.